OpenStudy (anonymous):

What is the value of p in the equation 4p + 3(p + 4) = 12p + 12? 12 1 0 −1

OpenStudy (anonymous):

A)12 B)1 C)0 D)-1

OpenStudy (anonymous):

please help

OpenStudy (kidrah69):

\[3(p+4)\] distribute

OpenStudy (kidrah69):

what do you get?

OpenStudy (anonymous):

well when I simfly the numbers i got none

OpenStudy (anonymous):

when i did a it got a big number like 612

OpenStudy (kidrah69):

ok thats wrong do what im telling you look let me write it out \[4p+3(p+4)=12p+12\]distribute \[3(p+4)\] get \[4p+3p+12=12p+12\] now we have to combine like terms on the left \[4p+3p=?\]

OpenStudy (anonymous):

-1

OpenStudy (kidrah69):

What is 4+3??

OpenStudy (kidrah69):

remember we are adding

OpenStudy (anonymous):

7

OpenStudy (kidrah69):

good so now we have \[7p+12=12p+12\] now subtract 12 from both sides they will cancel out so you will now have \[7p=12p+0\] subtract \[7p-12p=?\]

OpenStudy (anonymous):

-5

OpenStudy (kidrah69):

ok and now what is \[-5\div0=?\]

OpenStudy (anonymous):

0

OpenStudy (kidrah69):

So what is the final answer?

OpenStudy (anonymous):

C)0

OpenStudy (kidrah69):

good job! Welcome to openstudy! :)

OpenStudy (anonymous):

thank you